what is the resultant displacement if a man jogs 120m, east, and then walks 50m due south?


1
Expert's answer
2021-09-30T10:20:30-0400

The resultant displacement is


"d=\sqrt{120^2+50^2}=130\text{ m},\\\space\\
\theta=\arctan\frac{120}{50}=67.3°\text{ East of South}."